# Self-Organization and Spectral Mechanism of Attractor Landscapes in High-Capacity Kernel Hopfield Networks

### v5

@@
[…]
Self-Organization and Spectral Mechanism of Attractor Landscapes in High-Capacity Kernel Hopfield Networks
@@
{summary}
Kernel-based learning methods can dramatically increase the storage capacity of Hopfield networks, yet the dynamical mechanism behind this enhancement remains poorly understood. We address this gap by unifying the geometric analysis of the attractor landscape with the spectral theory of kernel machines. Using a novel metric, “Pinnacle Sharpness,” we first uncover a rich phase diagram of attractor stability, identifying a “Ridge of Optimization” where the network achieves maximal robustness under high-load conditions. Phenomenologically, this ridge is characterized by a “Force Antagonism,” where a strong driving force is balanced by a collective feedback force. Theoretically, we reveal that this phenomenon arises from a specific reorganization of the weight spectrum, which we term Spectral Concentration. Unlike a simple rank-1 collapse, our analysis shows that the network on the ridge self-organizes into a critical state: the leading eigenvalue is amplified to maximize global stability (Direct Force), while the trailing eigenvalues are preserved to maintain high memory capacity (Indirect Force). These findings provide a complete physical picture of how high-capacity associative memories are formed, demonstrating that optimal performance is achieved by tuning the system to a spectral “Goldilocks zone” between rank collapse and diffusion.
@@ 1 Introduction
    Theoretical Mechanism: We propose a spectral theory to explain these observations. We establish a link between the geometric stability and the leading eigenvalue of the kernel Gram matrix. Based on this link, we reveal that the “Ridge” corresponds to a state of Spectral Concentration. Unlike a simple rank-1 collapse which would destroy memory capacity, our analysis shows that the network on the ridge self-organizes into a critical state: the leading eigenvalue is amplified to maximize global stability, while the trailing eigenvalues are preserved to maintain the rank required for multi-pattern storage.
@@
    Unified Physical Picture: We demonstrate that the observed “Force Antagonism” is the dynamical manifestation of Spectral Concentration. The dominant spectral mode creates a deep potential funnel (Direct Force), while the subordinate modes manage inter-pattern interference (Indirect Force).
@@
The remainder of this paper is organized as follows. Section 2 reviews related work, positioning our study within the context of capacity theory and kernel associative memories. Section 3 introduces the network model and defines the geometric metrics used for our analysis. Section 4 presents the phenomenological results, detailing the discovery of the “Ridge of Optimization” and the associated “Force Antagonism.” Section 5 provides the theoretical core of this work, deriving the spectral mechanism of stability and validating the “Spectral Concentration” hypothesis. Finally, Section 6 discusses the broader implications of our findings, and Section 7 concludes the paper.
@@ 2.1 From Classical to Modern Hopfield Networks
The classical Hopfield network [1], governed by the Hebbian learning rule, is theoretically elegant but practically limited by a storage capacity of $P\approx 0.14N$ [2]. To overcome this, researchers have proposed various modifications to the energy function. Polynomial interaction terms were introduced to increase capacity tensorially [7, 8]. More recently, Krotov and Hopfield proposed Dense Associative Memories using rectified polynomial energy functions, significantly boosting capacity [3]. Demircigil et al. further extended this to exponential energy functions, achieving exponential storage capacity [9]. A pivotal development was the work by Ramsauer et al., which established the equivalence between these exponential Hopfield networks and the attention mechanism in Transformers [4]. These “Modern Hopfield Networks” achieve high capacity by designing steep energy landscapes (e.g., Log-Sum-Exp). In contrast, our approach retains the standard quadratic energy structure (in the feature space) but leverages a powerful learning algorithm (KLR) to sculpt the landscape adaptively. This offers a complementary perspective: high capacity can arise from learning dynamics rather than just architectural definition.
@@ 2.3 Geometry and Spectrum of Neural Dynamics
Understanding the geometry of neural dynamics has been a longstanding goal, often approached through statistical mechanics [14] or information geometry [15]. Amari and Wu pioneered the geometric analysis of kernel machines, demonstrating how a kernel function induces a Riemannian metric in the input space that governs the spatial resolution of classification [15]. Our work complements these theoretical foundations by introducing a data-driven spectral perspective. While statistical mechanics often assumes the thermodynamic limit ($N\to\infty$) with random patterns, and classical information geometry focuses on static classifiers, our spectral analysis of the “Ridge” and “Spectral Concentration” provides a concrete mechanism for the dynamics of finite-size networks. It explains how the system self-organizes its weight spectrum to balance stability (via the leading eigenvalue) and diversity (via the trailing eigenvalues), offering a bridge between abstract geometric theories and the actual stability of learned attractors.
@@ 3.2 Phenomenological Tool: Pinnacle Sharpness and Force Decomposition
We utilize the Pinnacle Sharpness, defined as the squared norm of the gradient of this function at a stored pattern:
@@ 4.1 The Phase Diagram of Attractor Stability
Figure 1: Phase diagram of attractor stability, quantified by the Pinnacle Sharpness $M$. The heatmap shows $\log_{10}M$ as a function of kernel locality $\gamma$ and storage load $P/N$. Three regimes are visible: (i) a local regime (large $\gamma$), (ii) a global, inefficient regime (small $\gamma$, low load), and (iii) a remarkable “Ridge of Optimization” (bright diagonal band) where the network achieves maximal stability. This ridge corresponds to the spectral “Goldilocks zone” identified in our theoretical analysis.
@@
    The “Ridge of Optimization”: The most remarkable feature of the phase diagram is the emergence of a “ridge” of extremely high Pinnacle Sharpness, extending diagonally from the lower-left to the lower-right. Along this ridge, the network achieves a level of attractor stability that is orders of magnitude greater than in the other regimes. This indicates that, faced with the dual challenges of a global kernel (small $\gamma$) and high memory congestion (large $P/N$), the network does not simply fail. Instead, it enters a highly optimized state, suggesting a sophisticated self-organization mechanism is at play.
@@ 4.2 Decomposing the Forces: The Mechanism of Self-Organization
Figure 2: Phase diagram of Force Interference $\rho$, measuring the cosine similarity between the Direct Force $\bm{F}_{d}$ and Indirect Force $\bm{F}_{i}$. Blue regions indicate strong anti-correlation (antagonism, $\rho\approx-1$). The “Ridge of Optimization” in Fig. 1 precisely traces the transition boundary into this antagonistic regime, suggesting that stability is maintained by a delicate balance of opposing spectral forces.
To understand the origin of the “Ridge of Optimization,” we analyzed the behavior of the two constituent forces of the landscape gradient: the direct force $\bm{F}_{d}$ and the indirect feedback force $\bm{F}_{i}$. We computed the normalized correlation between these two forces, the Force Interference $\rho(\bm{x})$, across the same parameter space.
@@ 4.3 The Nature of the Optimization Ridge
Finally, to elucidate the nature of the “Ridge of Optimization,” we analyzed the magnitudes of the individual forces along a representative cross-section of the phase diagram. We fixed gamma to a small value ($\gamma=0.001$), where the cooperative anti-correlation is strong, and plotted the log-magnitudes of $\|\bm{F}_{d}\|^{2}$, $\|\bm{F}_{i}\|^{2}$, and the resulting Pinnacle Sharpness ($\|\nabla V\|^{2}$) as a function of the storage load $P/N$.
@@
This leads to our final conclusion. The “Ridge of Optimization” is a regime where: (1) the direct and indirect forces are in a state of strong, cooperative anti-correlation, and (2) the learning algorithm, driven by the high storage load, amplifies the magnitude of the direct force $\bm{F}_{d}$ to a level that vastly overwhelms the opposing feedback force $\bm{F}_{i}$. The resulting large net force creates the exceptionally sharp and stable attractors observed. This is a remarkable demonstration of a system achieving extreme stability not through the synergy of forces, but through a controlled and amplified antagonism.
@@ 5 Theoretical Analysis: Spectral Mechanism of Self-Organization
The phenomenological experiments in Section 3 presented a compelling puzzle: the emergence of a “Ridge of Optimization” where stability is maximized amidst strong “Force Antagonism.” In this section, we solve this puzzle. We posit that these phenomena are manifestations of a spectral phase transition in the learned network. We first deduce this mechanism theoretically and then validate it with targeted experiments.
@@ 5.1 Theory: Spectral Mechanism of Stability
To provide a theoretical explanation for the observed “Ridge of Optimization” and the associated “Force Antagonism,” we analyze the spectral properties of the network in the feature space. We propose a two-step mechanism: first, we link the geometric stability to the spectral properties of the kernel, and second, we characterize the unique spectral state that emerges on the Ridge.
@@ 5.1.1 Step 1: The Geometry-Spectrum Link
The local stability of an attractor is geometrically quantified by the Pinnacle Sharpness $M$. Although the dynamics are defined by the kernel update rule, the underlying curvature of the energy landscape is governed by the Hessian matrix. By analyzing the Hessian in the reproducing kernel Hilbert space (see A for derivation), we establish a direct link between the attractor’s geometry and the spectrum of the learned model.
@@ 5.1.2 Step 2: Spectral Concentration on the Ridge
Based on Proposition 1, we can now interpret the “Ridge of Optimization.” We hypothesize that the high stability observed in this regime is not due to a uniform strengthening of all weights, but rather due to a specific reorganization of the spectral structure. We term this phenomenon Spectral Concentration.
@@
The “Force Antagonism” observed in Section 4 is the dynamical signature of this spectral hierarchy: the dominant mode exerts the primary restorative force, while the subordinate modes manage the interference to ensure separability.
@@ 5.2 Empirical Validation of the Theory
A lower Stable Rank indicates a more concentrated spectrum, where the leading mode dominates the energy. As shown in Fig. 4, the region where the Stable Rank is minimal perfectly overlaps with the “Ridge of Optimization” observed in Fig. 1. Note that while the rank is significantly reduced compared to the random initialization, it does not collapse to exactly 1.0, consistent with the requirement for multi-pattern storage.
Figure 4: Heatmap of the Stable Rank (effective rank) of the dual variable matrix $\bm{\alpha}$ across the phase diagram. Lower values (lighter/yellow colors) indicate a more concentrated spectrum. Crucially, the region of minimal Stable Rank perfectly aligns with the “Ridge of Optimization” observed in Fig. 1, confirming that maximal attractor stability is achieved when the weight spectrum becomes highly concentrated (but not fully collapsed to rank 1).
@@
  • On the Ridge: The spectrum (red solid line) exhibits a unique intermediate structure. It features a distinct leading mode ($\sigma_{1}$) that drives stability, yet crucially maintains a slowly decaying, heavy tail ($\sigma_{k>1}$). This confirms that the Ridge represents a “Goldilocks zone” where the network maximizes stability via spectral concentration without suffering from rank collapse.
These results provide direct empirical evidence that the “Ridge” corresponds to a phase where the network self-organizes into a state of extreme spectral concentration.
@@ 5.3 Unified Interpretation
The combination of phenomenological observation (Force Antagonism) and spectral evidence (Spectral Concentration) leads to a unified physical picture. The “Ridge” represents a sweet spot where the kernel locality $\gamma$ is tuned such that the system can amplify the principal spectral mode to form deep, sharp attractors (via Proposition 1), while effectively managing inter-pattern interference using the remaining spectral degrees of freedom. The “Force Antagonism” is the dynamical manifestation of this spectral disparity: the massive Direct Force arises from the dominant $\lambda_{1}$ component, while the Indirect Force reflects the collective contribution of the trailing modes, acting to refine the basin boundaries.
@@ 6 Discussion
In this work, we have moved beyond simple performance metrics to provide a clear physical picture of the self-organization mechanism in KLR-trained Hopfield networks. Our central finding is the discovery of the “Ridge of Optimization” and its spectral origin, “Spectral Concentration.” Here, we discuss the broader implications of these findings.
@@ 6.1 The Ridge as a Critical Phase Boundary
The spectral analysis in Fig. 5 revealed that the Ridge represents a critical state. The network does not simply maximize the leading eigenvalue to the point of rank-1 collapse, which would imply a trivial memory with unit capacity. Instead, it maintains a delicate balance: maximizing $\lambda_{1}$ for global stability while preserving a non-zero tail ($\lambda_{k>1}$) for pattern discriminability. This suggests that the Ridge can be interpreted as a phase boundary or a “Goldilocks zone” between an ordered phase (rank-1, low capacity) and a disordered phase (diffuse spectrum, weak stability). This observation aligns with the hypothesis of Self-Organized Criticality in biological neural networks, where systems tune themselves to a critical point to maximize information processing capabilities.
It is important to distinguish Spectral Concentration from simple dimensionality reduction techniques like PCA. PCA blindly discards trailing eigenvalues to minimize reconstruction error, often leading to the loss of pattern identity (similar to rank-1 collapse). In contrast, KLR learning is discriminative; it actively preserves the “heavy tail” of the spectrum exactly to the extent required to maintain the decision boundaries between patterns. Thus, the Ridge represents an optimal compression that balances energy minimization (via the leading mode) with classification margin (via the trailing modes).

### v6

@@
Abstract
Kernel-based learning methods can dramatically increase the storage capacity of Hopfield networks, yet the dynamical mechanism behind this enhancement remains poorly understood. We address this gap by unifying the geometric analysis of the attractor landscape with the spectral theory of kernel machines. Using a novel metric, Pinnacle Sharpness, we first uncover a rich phase diagram of attractor stability, identifying a Ridge of Optimization where the network achieves maximal robustness under high-load conditions. Phenomenologically, this ridge is characterized by a Force Antagonism, where a strong driving force is balanced by a collective feedback force. Theoretically, we reveal that this phenomenon arises from a specific reorganization of the weight spectrum, which we term Spectral Concentration. Unlike a simple rank-1 collapse, our analysis shows that the network on the ridge self-organizes into a critical state: the leading eigenvalue is amplified to maximize global stability (Direct Force), while the trailing eigenvalues are preserved to maintain high memory capacity (Indirect Force). These findings provide a complete physical picture of how high-capacity associative memories are formed, demonstrating that optimal performance is achieved by tuning the system to a spectral Goldilocks zone between rank collapse and diffusion.
@@ 1 Introduction
    Theoretical Mechanism: We propose a spectral theory to explain these observations. We establish a link between the geometric stability and the leading eigenvalue of the kernel Gram matrix. Based on this link, we reveal that the Ridge corresponds to a state of Spectral Concentration. Unlike a simple rank-1 collapse which would destroy memory capacity, our analysis shows that the network on the ridge self-organizes into a critical state: the leading eigenvalue is amplified to maximize global stability, while the trailing eigenvalues are preserved to maintain the rank required for multi-pattern storage.
@@
    Unified Physical Picture: We demonstrate that the observed Force Antagonism is the dynamical manifestation of Spectral Concentration. The dominant spectral mode creates a deep potential funnel (Direct Force), while the subordinate modes manage inter-pattern interference (Indirect Force).
@@
The remainder of this paper is organized as follows. Section 2 reviews related work, positioning our study within the context of capacity theory and kernel associative memories. Section 3 introduces the network model and defines the geometric metrics used for our analysis. Section 4 presents the phenomenological results, detailing the discovery of the Ridge and the associated Force Antagonism. Section 5 provides the theoretical core of this work, deriving the spectral mechanism of stability and validating the Spectral Concentration hypothesis. Finally, Section 6 discusses the broader implications of our findings, and Section 7 concludes the paper.
@@ 2.1 From Classical to Modern Hopfield Networks
The classical Hopfield network [1], governed by the Hebbian learning rule, is theoretically elegant but practically limited by a storage capacity of $P\approx 0.14N$ [2]. To overcome this, researchers have proposed various modifications to the energy function. Polynomial interaction terms were introduced to increase capacity tensorially [7, 8]. More recently, Krotov and Hopfield proposed Dense Associative Memories using rectified polynomial energy functions, significantly boosting capacity [3]. Demircigil et al. further extended this to exponential energy functions, achieving exponential storage capacity [9]. A pivotal development was the work by Ramsauer et al., which established the equivalence between these exponential Hopfield networks and the attention mechanism in Transformers [4]. These Modern Hopfield Networks achieve high capacity by designing steep energy landscapes (e.g., Log-Sum-Exp). In contrast, our approach retains the standard quadratic energy structure (in the feature space) but leverages a powerful learning algorithm (KLR) to sculpt the landscape adaptively. This offers a complementary perspective: high capacity can arise from learning dynamics rather than just architectural definition.
@@ 2.3 Geometry and Spectrum of Neural Dynamics
Understanding the geometry of neural dynamics has been a longstanding goal, often approached through statistical mechanics [14] or information geometry [15]. Amari and Wu pioneered the geometric analysis of kernel machines, demonstrating how a kernel function induces a Riemannian metric in the input space that governs the spatial resolution of classification [15]. Our work complements these theoretical foundations by introducing a data-driven spectral perspective. While statistical mechanics often assumes the thermodynamic limit ($N\to\infty$) with random patterns, and classical information geometry focuses on static classifiers, our spectral analysis of the Ridge and Spectral Concentration provides a concrete mechanism for the dynamics of finite-size networks. It explains how the system self-organizes its weight spectrum to balance stability (via the leading eigenvalue) and diversity (via the trailing eigenvalues), offering a bridge between abstract geometric theories and the actual stability of learned attractors.
@@ 3.2 Phenomenological Tool: Pinnacle Sharpness and Force Decomposition
We utilize the “Pinnacle Sharpness”, defined as the squared norm of the gradient of this function at a stored pattern:
@@ 4.1 The Phase Diagram of Attractor Stability
Figure 1: Phase diagram of attractor stability, quantified by the Pinnacle Sharpness $M$. The heatmap shows $\log_{10}M$ as a function of kernel locality $\gamma$ and storage load $P/N$. Three regimes are visible: (i) a local regime (large $\gamma$), (ii) a global, inefficient regime (small $\gamma$, low load), and (iii) a remarkable Ridge of Optimization (bright diagonal band) where the network achieves maximal stability. This ridge corresponds to the spectral “Goldilocks zone” identified in our theoretical analysis.
@@
    The Ridge of Optimization: The most remarkable feature of the phase diagram is the emergence of a “ridge” of extremely high Pinnacle Sharpness, extending diagonally from the lower-left to the lower-right. Along this ridge, the network achieves a level of attractor stability that is orders of magnitude greater than in the other regimes. This indicates that, faced with the dual challenges of a global kernel (small $\gamma$) and high memory congestion (large $P/N$), the network does not simply fail. Instead, it enters a highly optimized state, suggesting a sophisticated self-organization mechanism is at play.
@@ 4.2 Decomposing the Forces: The Mechanism of Self-Organization
Figure 2: Phase diagram of Force Interference $\rho$, measuring the cosine similarity between the Direct Force $\bm{F}_{d}$ and Indirect Force $\bm{F}_{i}$. Blue regions indicate strong anti-correlation (antagonism, $\rho\approx-1$). The Ridge of Optimization in Fig. 1 precisely traces the transition boundary into this antagonistic regime, suggesting that stability is maintained by a delicate balance of opposing spectral forces.
To understand the origin of the Ridge, we analyzed the behavior of the two constituent forces of the landscape gradient: the direct force $\bm{F}_{d}$ and the indirect feedback force $\bm{F}_{i}$. We computed the normalized correlation between these two forces, the Force Interference $\rho(\bm{x})$, across the same parameter space.
@@ 4.3 The Nature of the Optimization Ridge
Finally, to elucidate the nature of the Ridge, we analyzed the magnitudes of the individual forces along a representative cross-section of the phase diagram. We fixed gamma to a small value ($\gamma=0.001$), where the cooperative anti-correlation is strong, and plotted the log-magnitudes of $\|\bm{F}_{d}\|^{2}$, $\|\bm{F}_{i}\|^{2}$, and the resulting Pinnacle Sharpness ($\|\nabla V\|^{2}$) as a function of the storage load $P/N$.
@@
This leads to our final conclusion. The Ridge of Optimization is a regime where: (1) the direct and indirect forces are in a state of strong, cooperative anti-correlation, and (2) the learning algorithm, driven by the high storage load, amplifies the magnitude of the direct force $\bm{F}_{d}$ to a level that vastly overwhelms the opposing feedback force $\bm{F}_{i}$. The resulting large net force creates the exceptionally sharp and stable attractors observed. This is a remarkable demonstration of a system achieving extreme stability not through the synergy of forces, but through a controlled and amplified antagonism.
@@ 5 Theoretical Analysis: Spectral Mechanism of Self-Organization
The phenomenological experiments in Section 4 presented a compelling puzzle: the emergence of the Ridge where stability is maximized amidst strong Force Antagonism. In this section, we solve this puzzle. We posit that these phenomena are manifestations of a spectral phase transition in the learned network. We first deduce this mechanism theoretically and then validate it with targeted experiments.
@@ 5.1 Theory: Spectral Mechanism of Stability
To provide a theoretical explanation for the observed Ridge and the associated Force Antagonism, we analyze the spectral properties of the network in the feature space. We propose a two-step mechanism: first, we link the geometric stability to the spectral properties of the kernel, and second, we characterize the unique spectral state that emerges on the Ridge.
@@ 5.1.1 Step 1: The Geometry-Spectrum Link
The local stability of an attractor is geometrically quantified by the Pinnacle Sharpness $M$. Although the dynamics are defined by the kernel update rule, the underlying curvature of the energy landscape is governed by the Hessian matrix. By analyzing the Hessian in the reproducing kernel Hilbert space (see Appendix A.2 for derivation), we establish a direct link between the attractor’s geometry and the spectrum of the learned model.
@@ 5.1.2 Step 2: Spectral Concentration on the Ridge
Based on Proposition 1, we can now interpret the Ridge. We hypothesize that the high stability observed in this regime is not due to a uniform strengthening of all weights, but rather due to a specific reorganization of the spectral structure. We term this phenomenon Spectral Concentration.
@@
The Force Antagonism observed in Section 4 is the dynamical signature of this spectral hierarchy: the dominant mode exerts the primary restorative force, while the subordinate modes manage the interference to ensure separability.
@@ 5.2 Empirical Validation of the Theory
A lower Stable Rank indicates a more concentrated spectrum, where the leading mode dominates the energy. As shown in Fig. 4, the region where the Stable Rank is minimal perfectly overlaps with the Ridge observed in Fig. 1. Note that while the rank is significantly reduced compared to the random initialization, it does not collapse to exactly 1.0, consistent with the requirement for multi-pattern storage.
Figure 4: Heatmap of the Stable Rank (effective rank) of the dual variable matrix $\bm{\alpha}$ across the phase diagram. Lower values (lighter/yellow colors) indicate a more concentrated spectrum. Crucially, the region of minimal Stable Rank perfectly aligns with the Ridge of Optimization observed in Fig. 1, confirming that maximal attractor stability is achieved when the weight spectrum becomes highly concentrated (but not fully collapsed to rank 1).
@@
  • On the Ridge: The spectrum (red solid line) exhibits a unique intermediate structure. It features a distinct leading mode ($\sigma_{1}$) that drives stability, yet crucially maintains a slowly decaying, heavy tail ($\sigma_{k>1}$). This confirms that the Ridge represents a Goldilocks zone where the network maximizes stability via spectral concentration without suffering from rank collapse.
These results provide direct empirical evidence that the Ridge corresponds to a phase where the network self-organizes into a state of extreme spectral concentration.
@@ 5.3 Unified Interpretation
The combination of phenomenological observation (Force Antagonism) and spectral evidence (Spectral Concentration) leads to a unified physical picture. The Ridge represents a sweet spot where the kernel locality $\gamma$ is tuned such that the system can amplify the principal spectral mode to form deep, sharp attractors (via Proposition 1), while effectively managing inter-pattern interference using the remaining spectral degrees of freedom. The Force Antagonism is the dynamical manifestation of this spectral disparity: the massive Direct Force arises from the dominant $\lambda_{1}$ component, while the Indirect Force reflects the collective contribution of the trailing modes, acting to refine the basin boundaries.
@@ 6 Discussion
In this work, we have moved beyond simple performance metrics to provide a clear physical picture of the self-organization mechanism in KLR-trained Hopfield networks. Our central finding is the discovery of a Ridge of Optimization and its spectral origin, Spectral Concentration. Here, we discuss the broader implications of these findings.
@@ 6.1 The Ridge as a Critical Phase Boundary
The spectral analysis in Fig. 5 revealed that the Ridge represents a critical state. The network does not simply maximize the leading eigenvalue to the point of rank-1 collapse, which would imply a trivial memory with unit capacity. Instead, it maintains a delicate balance: maximizing $\lambda_{1}$ for global stability while preserving a non-zero tail ($\lambda_{k>1}$) for pattern discriminability. This suggests that the Ridge can be interpreted as a phase boundary or a Goldilocks zone between an ordered phase (rank-1, low capacity) and a disordered phase (diffuse spectrum, weak stability). This observation aligns with the hypothesis of Self-Organized Criticality in biological neural networks [16], where systems tune themselves to a critical point to maximize information processing capabilities.
It is important to distinguish Spectral Concentration from simple dimensionality reduction techniques like Principal Component Analysis (PCA). PCA blindly discards trailing eigenvalues to minimize reconstruction error, often leading to the loss of pattern identity (similar to rank-1 collapse). In contrast, KLR learning is discriminative; it actively preserves the “heavy tail” of the spectrum exactly to the extent required to maintain the decision boundaries between patterns. Thus, the Ridge represents an optimal compression that balances energy minimization (via the leading mode) with classification margin (via the trailing modes).